OpenStudy (anonymous):

At 12:00 noon, Isabel started walking east from her house along a staight road at 2 miles per hour. Her sister left the house to go running at 1:00 p.pm, heading south at 8 miles per hour, also along a staight road. Let the house be at the orgin of a coordinate system with north up. 1. Give coordinates of Isabel's position t hours after noon. Give the coordinates of her sister;s position at the same time. 2.Use the distance formula to write an expression for the distance between Isabel and her sister t hours after noon , assuming t>1. Simplify the expression.

jimthompson5910 (jim_thompson5910):

were you able to get anywhere?

OpenStudy (anonymous):

i probabliy could have gotten the second one if i knew how to do the first but not really

jimthompson5910 (jim_thompson5910):

ok a drawing may help

jimthompson5910 (jim_thompson5910):



OpenStudy (anonymous):

so is the coordinate (2t,0)

jimthompson5910 (jim_thompson5910):

Give coordinates of Isabel's position t hours after noon. well if t hours have gone by, then she is t*2 = 2t miles east of (0,0) so the coordinates of Isabel's position t hours after noon is (-2t, 0)

jimthompson5910 (jim_thompson5910):

oh wow...mixed up my east and west...really bad error...lol

jimthompson5910 (jim_thompson5910):

i feel dumb lol but yeah, it's (2t, 0)

OpenStudy (anonymous):

lmao its okay

OpenStudy (anonymous):

and her sisters is (8t,0) right?

jimthompson5910 (jim_thompson5910):

Give the coordinates of her sister;s position at the same time. her position is at (0, -8t) because she's going 8 mph south for t hours

OpenStudy (anonymous):

oh okay

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

for part b, we need to find the distance of that line segment

OpenStudy (anonymous):

okay so i set it up like d=\[\sqrt{(2t-(-8t))^2+(0-0))^2}\]

jimthompson5910 (jim_thompson5910):

no, but you have the right idea though

jimthompson5910 (jim_thompson5910):

it should be this \[\large d = \sqrt{(2t-0)^2+(0-8t))^2}\] then you simplify

OpenStudy (anonymous):

okay thanks soooo much lol

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

wait one more thing?

OpenStudy (anonymous):

u here?

jimthompson5910 (jim_thompson5910):

what's your question

OpenStudy (anonymous):

when it simplifies it is \[\sqrt{68t^2}\] right?

jimthompson5910 (jim_thompson5910):

you can go further

jimthompson5910 (jim_thompson5910):

\[\large \sqrt{68t^2}\] \[\large \sqrt{4t^2*17}\] \[\large \sqrt{4t^2}*\sqrt{17}\] \[\large 2t\sqrt{17}\]

jimthompson5910 (jim_thompson5910):

So the distance between them at time t is \[\large d = 2t\sqrt{17}\]

OpenStudy (anonymous):

then it says find the distance when it is 2 p.m.

jimthompson5910 (jim_thompson5910):

2 pm is 2 hours away from 12:00 pm noon